I have not really been to good at math and I never had a good Alg. teacher b/c they all quit or something happend so we were stuck w/diff. subs and they didn't make us do hw b/c they didn't understand Alg either. So question is for this Alg. 2 assign. deals w/equations of Lines
Going by 3's
there is a graph-- there is a pt on --going across -9 and down -6
then there is another pt. --going across -12 and down -18
it's asking to write an equation of the line. I don't know exactly how to interpret the graph b/c graphing for me is hard (to understand.)
P.S. My teacher said to use the y-y=m(x-x) formula but I am just totally confused please help me it would be greatly appreciated so I can do the other prob. I am having troubles w/.

Found 2 solutions by longjonsilver, richwmiller:
Answer by longjonsilver(2297)

please read my Lessons on linear algebra on this website.

I am a good teacher (honestly) and i believe what i have written is easy to follow. Have a go and see what you think.

Good luck

Jon.

Answer by richwmiller(17219)

y-y=m(x-x) will work just fine
Your points are (-12,-18) and (-9,-6)
y = 4x+30
-18+6=m(-12+9)
-12=-3m
-12/-3=m
y+6=4*(x+9)
y+6=4x+36
y=4x+30
or are you saying
(12,18) and (9,6)
your y's are 18 and 6
your x's are 12 and 9
18-6=m(12-9)
12=3m
12/3=m
4=m
y-6=4*(x-9)
y-6=4x-36
y=4x-30